I need 25 lb of a combination of product A 27 lb/cube ft & product B 13 lb/cube ft how much of each will I need. What is the formula?
Thank you

Answer by ankor@dixie-net.com(22740)

I need 25 lb of a combination of product A 27 lb/cube ft & product B 13 lb/cube ft how much of each will I need.
What is the formula?
:
You must specify the amt per cu/ft you want the resulting 25 lbs to be
It has to be between 27 and 13, as an example I am going to say 18 lb/cuft
:
Let a = amt of product A
then
(25-a) = amt of product B
:
27a + 13(25-a) = 18(25)
27a + 325 - 13a = 450
27a - 13a = 450 - 325
14a = 125
a = 125/14
a = 8.93 lb of product a
then
(25-8.93) = 16.07 lb of product b
:
:
If you wish to have a resulting 25 lb which 20 lb/cuft, half way between 13 & 27
just have 12.5 lb of each product